Letter Z is 26 in base ten;decide Z in binary scale
Found 2 solutions by Solver92311, greenestamps:
Answer by Solver92311(821)











So




John

My calculator said it, I believe it, that settles it

From
I > Ø

Answer by greenestamps(13200)



In general, to convert a number n in base 10 to another base b, repeatedly divide n by b and pick off the remainders, until the number left is 0.

To convert 26 base 10 to base 2...

26/2=13 remainder 0
13/2 = 6 remainder 1
6/2 = 3 remainder 0
3/2 = 1 remainder 1
1/2 = 0 remainder 1

Those remainders, written in reverse order, are the digits of the converted number in base 2.

ANSWER: 26 (base 10) = 11010 (base 2)

Here is another random example to help you understand the process: convert 417 base 10 to base 5:

417/5 = 83 remainder 2
83/5 = 16 remainder 3
16/5 = 3 remainder 1
3/5 = 0 remainder 3

ANSWER: 417 (base 10) = 3132 (base 5)

CHECK: 3132 (base 5) = 3(5^3)+1(5^2)+3(5^1)+2(5^0)=3(125)+1(25)+3(5)+2=375+25+15+2=417